### v1

@@ Abstract
Context:Calibration of optical polarimeters relies on the use of stars with negligible polarization (unpolarized standard stars) for determining the instrumental polarization zero-point. For wide-field polarimeters, calibration is often done by imaging the same star over multiple positions in the field of view - a process which is time-consuming. A more effective technique is to target fields containing multiple standard stars. While this method has been used for fields with highly polarized stars, there are no such sky regions with well-measured unpolarized standard stars.
Aims:We aim to identify sky regions with tens of stars exhibiting negligible polarization, which are suitable for zero-point calibration of wide-field polarimeters.
Methods:We selected stars in regions with extremely low reddening, located at high Galactic latitudes. We targeted four fields in the northern, and eight in the southern Equatorial hemisphere. Observations were carried out at the Skinakas Observatory and the South African Astronomical Observatory respectively.
Results:We find two fields in the North and seven in the South with mean polarization lower than .
Conclusions:At least nine out of twelve fields can be used for zero-point calibration of wide-field polarimeters.
Zero-polarization candidate regions for calibration of wide-field optical polarimeters
@@ 1 Introduction
To date, the development of optical polarimetry instrumentation has primarily emphasized the creation of instruments tailored for single-source observations or with limited field-of-view capabilities. The calibration process for these instruments typically involves utilizing polarimetric standard stars (e.g. Blinov et al., 2023). When there is a need to calibrate the entire field of view, calibration methods often employ techniques such as rastering, where a star of known polarization is placed in different positions of the CCD to assess the instrument-induced polarization variations accross the field (e.g. King et al., 2014). This can be very time-consuming and, therefore inefficient for wide field instruments with lots of CCD pixels, especially in the scheme of wide-field surveys.
@@
The domain of optical polarimetry is on the verge of a significant transition as it enters the era of large-area surveys like Pasiphae (Tassis et al., 2018), SouthPol (Magalhães et al., 2012), and VSTPol (Covino et al., 2020). These surveys will utilize polarimeters equipped with unprecedentedly wide fields of view. However, the calibration process for these wide-field polarimeters presents a formidable challenge, particularly considering the surveys’ aim of achieving high accuracy levels of to fulfill their scientific objectives.
@@
Pasiphae is a pioneering survey with the primary objective of mapping the polarization of millions of stars located away from the Galactic plane. The scientific goals of the project include the determination of dust cloud distributions and their properties along each line of sight, as well as the investigation of the magnetic field structure within our Galaxy (Pelgrims et al., 2023). Its findings will facilitate the accurate removal of the dust polarization foreground for surveys targeting the detection of the imprint of the primordial polarization B-modes on the cosmic microwave background. (e.g. Polnarev, 1985; Kamionkowski et al., 1997). Pasiphae will utilize two Wide-Area Linear Optical Polarimeters (WALOPs) which will operate from both the southern and northern hemispheres and will be mounted on the South African Astronomical Observatory’s (SAAO) 1 m telescope111https://www.saao.ac.za/, and Skinakas Observatory 1.3 m telescope222https://skinakas.physics.uoc.gr/, respectively. The description of the optical design, expected performance, and overall instrument design can be found in Maharana et al. (2020, 2021).
@@
Given that the targeted accuracy of Pasiphae in polarization fraction is , it is important to calibrate the whole field of view of WALOPs down to this level. Maharana et al. (2022) characterized the magnitude and sources of instrumental polarization and developed an on-sky polarimetric calibration method to obtain the target accuracy for WALOP-South. The calibration model creates an accurate mapping function between the instrument measured, and real Stokes parameters of a target. To develop the calibration model, wide-field sources of known and constant (also referred to as ”flat”) polarization across the field are observed through the instrument. The corresponding measured polarization values are used to create the mapping functions. One such calibration source is the sky on bright-Moon nights in the vicinity of the Moon. Patat & Romaniello (2006) calibrated the VLT-FORS1 instrument, using the qualitative assumption that the Moon-induced night sky polarization remains fairly constant in scales of a few arcminutes, while González-Gaitán et al. (2020) calibrated the FORS2 instrument by explicitly calculating the polarization of the scattered Moonlight. Maharana et al. (2023) have demonstrated that the sky polarization remains constant at the level of or less for fields of size 10 to 20 arcminutes up to 20 degrees away from the Moon, making them suitable ”polarimetric flat fields”. As shown by Maharana et al. (2022), the polarimetric flat fields suffice for calibration of wide field polarimeters, when used in conjugation with polarimetric standard stars. While polarimetric flat sources allow for relative calibration of the whole field, standard star measurements allow the estimation of the true polarization of the polarimetric flat source, enabling complete and absolute calibration of the entire field of view.
@@
Star fields with uniform polarization across the field can serve as candidate on-sky wide-field calibrators, provided their polarization scatter is smaller than the required calibration accuracy. The advantages of this compared to the full-Moon sky are that (1) it can be carried out on any given night, provided that such a field is available, and (2) it can be used as standard calibrator source whose polarization is known directly from observations, unlike the full-Moon sky, whose polarization can only be calculated based on modelling, or needs multiple observations in different angles from the moon (González-Gaitán et al., 2020) to achieve a similar result to an observation of a zero-polarized field. This eliminates the need for additional standard star measurements needed to estimate the true polarization of the polarimetric flat source. A major, and often the dominant result of instrumental polarization that wide-field polarimeters suffer from is the polarimetric zero offsets, i.e., the measured Stokes parameters when the observed source is unpolarized. This needs to be corrected using unpolarized wide-field targets. By observing unpolarized targets, one can directly get the polarimetric zero offset, as any potential non-zero cross-talk terms between the Stokes and are eliminated. We note that in polarimeters that show non-negligible cross-talk between the linear and circular Stokes parameters, zero-polarization standards are not enough to address this issue. Calibration in these cases would need more extensive modelling (Wiersema et al., 2018), observatoins of Stokes polarization standards (Giro et al., 2003), and/or in-built calibration sources/devices within the instrument, like in the case if WALOP (Maharana et al., 2022).
@@
Nonethelss, observations of zero-polarization fields are always essential. To fulfill this missing ingredient for calibrating wide-field polarimeters, in this work we aim to search for zero-polarized patches of the sky that can serve as calibrators for the WALOPs and other wide-field polarimeters. Skalidis et al. (2018) surveyed three regions of the northern sky at high Galactic latitudes with RoboPol, and found one to be consistent with zero polarization, while Clemens et al. (2012) used polarized stellar cluster observations to calibrate the Mimir near-infrared imaging wide-field polarimeter (Clemens et al., 2007). In order to search for zero-polarized patches, we rely on the known connection between extinction by interstellar dust and observed polarization in a given line-of-sight.
@@
Interstellar dust is ubiquitous in all lines-of-sight (e.g. Planck Collaboration et al., 2014). Its interaction with the interstellar magnetic field makes interstellar dust clouds act as polarizing filters. Asymmetric interstellar dust grains tend to align their minor axis parallel to the magnetic field lines (Andersson et al., 2015). Dichroic absorption in optical wavelengths (i.e. the attenuation of the component of the light polarized in the plane-of-the-sky direction of the major axis of the grains) gives rise to dust-induced polarization to originally unpolarized sources behind said dust, with its orientation along the minor axis of the grains, i.e. along the magnetic field lines. In the infrared and sub-mm regime, dust emits thermally, with the emission being polarized along the direction of its major axis, which is perpendicular to the polarization direction observed in the optical. The polarized emission from dust grains has been recently mapped with Planck and shown to be well-correlated with the optical starlight polarization (Planck Collaboration et al., 2020a).
@@
In addition to dust-induced polarization, light passing through the dust suffers from reddening as well. The magnitude of these effects is related to the amount of dust along the line of sight. Empirical investigations have established a relationship between the maximum observed polarization and the reddening , expressed as (Panopoulou et al., 2019; Planck Collaboration et al., 2020a). In this work we use this empirical relation, in conjunction with a publicly available reddening map to identify regions with expected polarization , that could be used as calibrators for wide-field polarimeters. We conducted polarimetric observations in the optical to verify whether these regions are indeed negligibly polarized. In Sect. 2 we describe the sample selection, data acquisition and processing. In Sect. 3 we present the results of the polarization observations. We make our conclusions in Sect. 4, and provide supplementary plots in Appendices A and B.
@@ 2.1 Sample selection
Our goal is to target regions in the sky with minimal polarization and with an appreciable number of stars that can be used for determining the instrumental zero-point across a field of view of , which is slightly wider than the field of view of the WALOPs that will be used for Pasiphae. We are not interested to only find specific zero-polarized stars in wide fields, such as fields with globular or open clusters, (e.g, Clemens et al., 2012). Such fields may indeed host clusters that are close enough so that their polarization is not affected by the interstellar medium (ISM). However, in these fields there may exist stars outside the cluster that are farther away and well-affected by the ISM. Rather, asserting that a certain region’s polarization is minimally affected, ensures that any given star in this region, regardless of distance, or brightness will be suitable for calibration, except for cases of intrinsically polarized stars, which can be easily identified. Thus, the calibration procedure becomes simpler, and stars used for calibration do not necessarily need to be exactly the ones used in this specific work. For example, the use of other, potentially fainter, sources in our targeted regions would be equally qualified as calibrators, thus making these regions useful for bigger telescopes as well, where bright nearby targets would saturate their cameras.
@@
Existing optical polarization data are sparse at high Galactic latitudes, with a mean density of 0.1 measurements per square degree (Panopoulou et al., 2023). One avenue would be to identify regions with minimal polarized dust emission as observed by Planck. However, the Planck polarization data are dominated by noise at high Galactic latitudes (Planck Collaboration et al., 2020c, b). Therefore, we searched for the lowest-polarization regions of the sky through reddening, which is an indirect observable of expected polarization levels, as discussed in Sect. 1. We used the Lenz et al. (2017) reddening map, as it is considered the most accurate at high Galactic latitudes (Chiang & Ménard, 2019). Based on this map, we selected regions with mag, corresponding to a maximum expected polarization fraction of , according to the relation (Panopoulou et al., 2019; Planck Collaboration et al., 2020a). This resulted to 799 square degrees of usable sky area, compared to the total of 16,342 square degrees in the map.
@@
Within the selected area, we searched for regions which fulfilled the maximum reddening of 0.01 mag criterion across the entire region. We aimed to find fields as spread-out in Right Ascension (RA) as possible, within the visibility constraints of the locations of the Pasiphae telescopes. This was to ensure that at least one such region is available on any observing night for calibration purposes. However, such regions are not available in a wide range of RAs, as we would wish for. Only 4 regions at northern latitudes and 8 regions at southern latitudes satisfied the aforementioned criteria, which we surveyed. Hereafter, we refer to these regions as ”dark patches” (DPN for North, DPS for South).
@@
Within each region, we selected all stars in Gaia DR2 (Gaia Collaboration et al., 2016, 2018) (which was the latest available dataset at the time of target selection) with band magnitude mag, to observe as many stars as possible in the least amount of time. Stars that were too bright ( mag) were excluded to avoid saturation on the CCD. We further excluded stars marked as variable in the Gaia dataset. Stars with overlapping point-spread-functions as seen in Digital Sky Survey (DSS)333https://archive.eso.org/dss/dss images of the fields were also excluded, as their polarization could be a result of spurious measurements.
2.2 Data aquisition and reduction - North
@@
In order to acquire data for the dark patch candidates in the northern hemisphere, we utilized the RoboPol instrument, mounted on the Skinakas 1.3 m telescope in Crete, Greece. RoboPol contains no rotating parts and can measure the linear Stokes parameters and with a single exposure. Its novel design allows for minimization of systematic and random errors (Ramaprakash et al., 2019). The instrument is optimized for single target measurements in the center of its field of view, where the background noise is minimized with the use of a mask.
@@
We observed four candidate patches between May and November 2021 and June to August 2023. The observations were conducted in Sloan Digital Sky Survey (SDSS)-r and Johnsons-Cousins R bands. Observations of both bands were considered together, as we have verified with standard stars measurements that these filters are essentially equivalent (Blinov et al., 2023). We selected these filters as the upcoming Pasiphae survey will be operating in SDSS-r. Data reduction and calibration were performed with the standard RoboPol pipeline as outlined in King et al. (2014); Panopoulou et al. (2015); Blinov et al. (2021). Uncertainties of the systematic error are of the order of in the mask of the instrument (Blinov et al., 2023). We note that these uncertainties concern measurements of individual targets. However, when we study a particular field, the accuracy to which we know its mean polarization is dependent on the number of stars observed in that field. The uncertainty of the mean polarization of the field decreases with the number of stars (refer to Eq. 5), and can ultimately be lower than the uncertainty in the polarization of individual targets.
2.3 Data acquisition and reduction - South

### v2

@@ Abstract
Context. The calibration of optical polarimeters relies on the use of stars with negligible polarization (i.e., unpolarized standard stars) for determining the instrumental polarization zero point. For wide-field polarimeters, calibration is often done by imaging the same star over multiple positions in the field of view (FoV), which is a time-consuming process. A more effective technique is to target fields containing multiple standard stars. While this method has been used for fields with highly polarized stars, there are no such sky regions with well measured unpolarized standard stars.
Aims. We aim to identify sky regions with tens of stars exhibiting negligible polarization that are suitable for a zero-point calibration of wide-field polarimeters.
Methods. We selected stars in regions with extremely low reddening, located at high Galactic latitudes. We targeted four fields in the northern and eight in the southern equatorial hemispheres. Observations were carried out at the Skinakas Observatory and the South African Astronomical Observatory.
Results. We found two fields in the north and seven in the south characterized by a mean polarization lower than .
Conclusions. At least 9 out of the 12 fields can be used for a zero-point calibration of wide-field polarimeters.
@@ 1 Introduction
To date, the development of optical polarimetry instrumentation has primarily relied on the creation of instruments tailored for single-source observations or with limited field-of-view (FoV) capabilities. The calibration process for these instruments typically involves utilizing polarimetric standard stars (e.g., Blinov et al., 2023). When there is a need for calibrating the entire FoV, calibration methods often employ techniques such as rastering, where a star of known polarization is placed in different positions of the CCD to assess the instrument-induced polarization variations accross the field (e.g., King et al., 2014). This can be very time-consuming and, thus, inefficient for wide field instruments with a great deal of CCD pixels, especially in the scheme of wide-field surveys.
@@
The domain of optical polarimetry is on the verge of a significant transition as it enters the era of large-area surveys such as Pasiphae (Tassis et al., 2018), SouthPol (Magalhães et al., 2012), and VSTPol (Covino et al., 2020). These surveys will utilize polarimeters equipped with unprecedentedly wide FoVs. However, the calibration process for these wide-field polarimeters presents a formidable challenge, particularly considering the surveys’ aim of achieving high accuracy levels of to fulfill their scientific objectives.
@@
Pasiphae is a pioneering survey with the primary objective of mapping the polarization of millions of stars located away from the Galactic plane. The scientific goals of the project include the determination of dust cloud distributions and their properties along each line of sight (LoS), as well as the investigation of the magnetic field structure within our Galaxy (Pelgrims et al., 2023). Its findings will facilitate the accurate removal of the dust polarization foreground for surveys targeting the detection of the imprint of the primordial polarization B-modes on the cosmic microwave background. (e.g., Polnarev, 1985; Kamionkowski et al., 1997). Pasiphae will utilize two Wide-Area Linear Optical Polarimeters (WALOPs) which will operate from both the southern and northern hemispheres, mounted on the South African Astronomical Observatory’s (SAAO) 1 m telescope111https://www.saao.ac.za/ and Skinakas Observatory 1.3 m telescope222https://skinakas.physics.uoc.gr/, respectively. The description of the optical design, expected performance, and overall instrument design can be found in Maharana et al. (2020, 2021).
@@
Given that the targeted accuracy of Pasiphae in polarization fraction is , it is important to calibrate the whole FoV of WALOPs down to this level. Maharana et al. (2022) characterized the magnitude and sources of instrumental polarization and developed an on-sky polarimetric calibration method to obtain the target accuracy for WALOP-South. The calibration model creates an accurate mapping function between the instrument measured and the real Stokes parameters of a target. To develop the calibration model, wide-field sources of known and constant (also referred to as ”flat”) polarization across the field are observed through the instrument. The corresponding measured polarization values are used to create the mapping functions. One such calibration source is the sky on bright-Moon nights in the vicinity of the Moon. Patat & Romaniello (2006) calibrated the VLT-FORS1 instrument, using the qualitative assumption that the Moon-induced night sky polarization remains fairly constant in scales of a few arcminutes, while González-Gaitán et al. (2020) calibrated the FORS2 instrument by explicitly calculating the polarization of scattered moonlight. Maharana et al. (2023) have demonstrated that the sky polarization remains constant at the level of or less for fields of size 10 to 20 arcminutes up to 20 degrees away from the Moon, making them suitable ”polarimetric flat fields.” As shown by Maharana et al. (2022), the polarimetric flat fields suffice for calibration of wide field polarimeters, when used in conjugation with polarimetric standard stars. While polarimetric flat sources allow for relative calibration of the whole field, standard star measurements allow the estimation of the true polarization of the polarimetric flat source, enabling a complete and absolute calibration of the entire FoV.
@@
Star fields with uniform polarization across the field can serve as candidate on-sky wide-field calibrators, provided their polarization scatter is smaller than the required calibration accuracy. The advantages of this compared to the full Moon sky are: (1) it can be carried out on any given night, provided that such a field is available and (2) it can be used as standard calibrator source whose polarization is known directly from observations; this is unlike the full-Moon sky, whose polarization can only be calculated based on modeling or requires multiple observations at different angles from the moon (González-Gaitán et al., 2020) to achieve a similar result to an observation of a zero-polarized field. This novel approach eliminates the need for additional standard star measurements needed to estimate the true polarization of the polarimetric flat source. A major (and often dominant) result of instrumental polarization that wide-field polarimeters tend to suffer from is the polarimetric zero offsets, namely: the measured Stokes parameters when the observed source is unpolarized. This needs to be corrected using unpolarized wide-field targets. By observing unpolarized targets, we can directly obtain the polarimetric zero offset, as any potential non-zero cross-talk terms between the Stokes and are eliminated. We note that in polarimeters that show non-negligible cross-talk between the linear and circular Stokes parameters, zero-polarization standards are not enough to address this issue. In such cases, the calibration would require more extensive modeling (Wiersema et al., 2018), observations of Stokes polarization standards (Giro et al., 2003) and/or in-built calibration sources and/or devices within the instrument, as in the case of WALOP (Maharana et al., 2022).
@@
Nonetheless, observations of zero-polarization fields are always essential. To fill in this missing ingredient for calibrating wide-field polarimeters, here we aim to search for zero-polarized patches of the sky that can serve as calibrators for the WALOPs and other wide-field polarimeters. Skalidis et al. (2018) surveyed three regions of the northern sky at high Galactic latitudes with RoboPol and found one to be consistent with zero polarization, while Clemens et al. (2012) used polarized stellar cluster observations to calibrate the Mimir near-infrared imaging wide-field polarimeter (Clemens et al., 2007). In order to search for zero-polarized patches, we rely on the known connection between extinction by interstellar dust and observed polarization in a given LoS.
@@
Interstellar dust is ubiquitous in every LoS (e.g., Planck Collaboration et al., 2014). Its interaction with the interstellar magnetic field makes interstellar dust clouds act as polarizing filters. Asymmetric interstellar dust grains tend to align their minor axis parallel to the magnetic field lines (Andersson et al., 2015). Dichroic absorption in optical wavelengths (i.e., the attenuation of the component of the light polarized in the plane-of-the-sky direction of the major axis of the grains) gives rise to dust-induced polarization to originally unpolarized sources behind said dust, with its orientation along the minor axis of the grains (i.e., along the magnetic field lines). In the infrared (IR) and sub-mm regime, dust emits light thermally, with the emission being polarized along the direction of its major axis, which is perpendicular to the polarization direction observed in the optical. The polarized emission from dust grains has been recently mapped with Planck and shown to be well correlated with the optical starlight polarization (Planck Collaboration et al., 2020a).
@@
In addition to dust-induced polarization, light passing through the dust suffers from reddening as well. The magnitude of these effects is related to the amount of dust along the LoS. Empirical investigations have established a relationship between the maximum observed polarization and the reddening , expressed as (Panopoulou et al., 2019; Planck Collaboration et al., 2020a). In this work we use this empirical relation, in conjunction with a publicly available reddening map to identify regions with expected polarization, , that could be used as calibrators for wide-field polarimeters. We conducted polarimetric observations in the optical to verify whether these regions are indeed negligibly polarized. In Sect. 2, we describe the sample selection, data acquisition, and processing. In Sect. 3, we present the results of the polarization observations. We give our conclusions in Sect. 4 and provide supplementary plots in Appendices A and B.
@@ 2.1 Sample selection
Our goal is to target regions in the sky with minimal polarization and with an appreciable number of stars that can be used for determining the instrumental zero point across a FoV of , which is slightly wider than the FoV of the WALOPs that will be used for Pasiphae. We are not only interested in finding specific zero-polarized stars in wide fields, such as fields with globular or open clusters, (e.g., Clemens et al., 2012). Such fields may indeed host clusters that are close enough so that their polarization is not affected by the interstellar medium (ISM). However, these fields may feature stars outside the cluster that are farther away and fully affected by the ISM. Thus, asserting that a certain region’s polarization is minimally affected ensures that any given star in this region, regardless of distance, or brightness will be suitable for calibration, except in cases of intrinsically polarized stars, which can be easily identified. Thus, the calibration procedure becomes simpler, and the stars used for calibration do not necessarily need to be exactly the ones used in this specific work. For example, the use of other, potentially fainter, sources in our targeted regions would be equally qualified as calibrators, thus making these regions useful for bigger telescopes as well, where bright nearby targets would saturate their cameras.
@@
Existing optical polarization data are sparse at high Galactic latitudes, with a mean density of 0.1 measurements per square degree (Panopoulou et al., 2023). One avenue would be to identify regions with minimal polarized dust emission as observed by Planck. However, the Planck polarization data are dominated by noise at high Galactic latitudes (Planck Collaboration et al., 2020c, b). Therefore, we searched for the lowest-polarization regions of the sky through reddening, which is an indirect observable of the expected polarization levels, as discussed in Sect. 1. We used the Lenz et al. (2017) reddening map, as it is considered the most accurate at high Galactic latitudes (Chiang & Ménard, 2019). Based on this map, we selected regions with mag, corresponding to a maximum expected polarization fraction of , according to the relation (Panopoulou et al., 2019; Planck Collaboration et al., 2020a). This resulted to 799 square degrees of usable sky area, compared to the total of 16,342 square degrees in the map.
@@
Within the selected area, we searched for regions that fulfilled the maximum reddening of 0.01 mag criterion across the entire region. We aimed to find fields as spread-out with respect to the right ascension (RA) as possible, within the visibility constraints of the locations of the Pasiphae telescopes. This was to ensure that at least one such region is available on any observing night for calibration purposes. However, such regions are not available for as wide a range of RAs as we would wish for. Only four regions at northern latitudes and eight regions at southern latitudes available to our study satisfy the aforementioned criteria. Hereafter, we refer to these regions as ”dark patches” (DPN for north, DPS for south).
@@
Within each region, we selected all stars in Gaia DR2 (Gaia Collaboration et al., 2016, 2018), which was the latest available dataset at the time of target selection, with band magnitude mag, to observe as many stars as possible in the least amount of time. Stars that were too bright ( mag) were excluded to avoid saturation on the CCD. We further excluded stars marked as variable in the Gaia dataset. Stars with overlapping point spread functions, as seen in Digital Sky Survey (DSS)333https://archive.eso.org/dss/dss images of the fields, were also excluded, as their polarization could be a result of spurious measurements.
2.2 Data acquisition and reduction: North
@@
In order to acquire data for the dark patch candidates in the northern hemisphere, we utilized the RoboPol instrument, mounted on the Skinakas 1.3 m telescope in Crete, Greece. RoboPol contains no rotating parts and can measure the linear Stokes parameters and with a single exposure. Its novel design allows for systematic and random errors to be minimized (Ramaprakash et al., 2019). The instrument is optimized for single target measurements in the center of its FoV, where the background noise is minimized with the use of a mask.
@@
We observed four candidate patches between May and November 2021 and June to August 2023. The observations were conducted in Sloan Digital Sky Survey (SDSS)-r′ and Johnsons-Cousins R bands. Observations of both bands were considered together, as we verified with standard stars measurements that these filters are essentially equivalent (Blinov et al., 2023). The selection of these filters was also informed by the fact that the upcoming Pasiphae survey will be operating in SDSS-r′. The data reduction and calibration were performed with the standard RoboPol pipeline, as outlined in King et al. (2014); Panopoulou et al. (2015); Blinov et al. (2021). Uncertainties of the systematic error are on the order of in the mask of the instrument (Blinov et al., 2023). We note that these uncertainties concern measurements of individual targets. However, when we studied a particular field, the accuracy with of its mean polarization measurement is dependent on the number of stars observed in that field. The uncertainty of the mean polarization of the field decreases with the number of stars (refer to Eq. 5) and may ultimately be lower than the uncertainty in the polarization of individual targets.
2.3 Data acquisition and reduction: South